The SS Hindenburg

contained about 5.33 × 105 g of H2 gas when it burned at Lakehurst, New Jersey, in 1937. The chemical reaction is as follows:


2H2 + O2→ 2H2O


How many grams of H2O were produced?



1
Expert's answer
2020-09-21T06:41:39-0400

M(H2) = 2 g/mol

M(H2O) = 18 g/mol

n=m/M

n(H2)="\frac{5.33\times10^5}{2}" = 2.665×105 mol

Proportion from the reaction:

n(H2) = n(H2O) = 2.665×105 mol

m(H2O)= 18×2.665×105 = 47.97×105 g

Answer: 47.97×105 g